Current CO2 composition in air is about 380 ppm (mol/mol). What is [CO2 (aq)] dissolved in pure water, given Henry’s constant 0.034 M/atm. Assume air pressure is 1.00 atm.

I got 1.29x10-5 M when I did my calculations.
Answer #1

The concentration (C) = K * P, where K is Henry's constant and P is partial pressure of CO2

Partial pressure = mole fraction * Pressure = 380 * 10-6 * 1 = 3.8 * 10-4 atm

C = 0.034 * 3.8 * 10-4 = 1.29 * 10-5 M, your answer is right :)